Consider triangle ABC in the plane where
A=(-1,15)
B=(-6,8)
C=(-27,23)


Find the lengths of the sides of the triangle:
AB=
BC=
AC=

Answer by checkley77(12569)

A=(-1,15)
B=(-6,8)
C=(-27,23)
AB^2=(6-1)^2+(15-8)^2=5^2+7^2=25+49=74
AB=SQRT74=8.6 ANSWER.
BC^2=(27-6)^2+(23-8)^2=21^2+15^2=441+225=666
BC=SQRT666=25.8 ANSWER.
AC^2=(27-1)^2+(23-15)^2=26^2+8^2=676+64=740
AC=SQRT740=27.2 ANSWER.